Solve by elimination
2x + 3y = 21
5x - 2y = -14

Answer by praseena(37)

2x+3y=21
5x-2y=-14
Mltiply the first equation by 2 ==> 2 (x+3y) = 2*21
==> 2*x+2*3y=42
==>4x+6y=42............(a)
multiply second equation by 3 ==>3(5x-2y)=3*-14
==>15x-6y=-42 .......(b)
Now add equations (a) and (b)
4x+15x+6y+-6y=42+-42
19x=0
==>x=0
now substitute the value x=0 in any equation given in the question.
==>2*0+3*y=21
3*y=21
y=21/3 = 7

so the solution is x=0 and y=7